# On the convergence of bootstrap current to the Shaing-Callen limit in stellarators

### v1

@@ 2.3 Finite collisionality, numerical examples
An even part of the distribution function perturbation $g_{\text{even}}=g^{\text{even}}$ responsible for particle flux (Ware pinch) in the adjoint problem is shown in Fig. 6 for the same cases as in Fig. 3. As expected, boundary layer appears at low collisionality only at the trapped-passing boundary. The structure of this layer well agrees with the complete analytical solution (51), i.e. $g_{\text{even}}$ is an anti-symmetric function of $\varphi$ with respect to the middle stellarator symmetry point $\varphi_{b}$.
Similar to the direct problem, if the trapped-passing boundary layer is well separated from the nearest class transition boundary (lowest collisionality case, $\nu_{\ast}=10^{-6}$) the Ware pinch coefficient agrees with the asymptotic limit. At higher collisionalities where this boundary layer enters a class-transition boundary, one can observe increased $g_{\text{even}}$ in the local ripple wells whose lowest maximum determines this transition boundary.
@@
Figure 7: Even part of the distribution function $g_{\text{even}}$ as function of $\eta$ in the middle of the left off-set well, $\varphi=4\pi/3$ for various collisionalities (left) and $\iota=1/4$. The trapped-passing boundary is shown by a vertical line. The value of $g_{\text{even}}$ at the off-set well bottom, $g_{\text{bot}}$, is shown as a function of $\nu_{\ast}$ in the right plot.
This accumulation of the particles in local ripples drives the off-set of Ware pinch since $g_{\text{even}}$ there is anti-symmetric with respect to mid-point $\varphi_{b}=5\pi$, what is true also for geodesic curvature and, respectively, for bounce-averaged radial drift velocity in these local ripples. Therefore, particles in both ripples drive the flux of the same sign. In the following, we will call such ripples “off-set wells”. As one can see from Fig. 7, $g_{\text{even}}$ is constant in most of the off-set well, and changes only in the vicinity of class transition boundary. Therefore, we can characterize $g_{\text{even}}$ in the whole off-set well by its value at the off-set well bottom, which will be denoted as $g_{\text{off}}$. It can be seen that $g_{\text{off}}$ has similar feature to $\lambda_{bB}$, i.e. it changes sign with reducing collisionality before going to zero.
Properties of this distribution function off-set, which was also observed in GSRAKE modelling (Beidler, 2020), will be studied in more detail in the next section. Here, we only show an extra example in Fig. 8
@@
where more than one type of off-set wells is present simultaneously with different types of off-sets dominating at different collisionalities. In this example, both, off-set wells containing only the lowest class of trapped particles who traverse a single minimum during their bounce period (as in Fig. 6) and an off-set wells containing many classes are shown. Note that trapped-passing boundary layer which has a standard (collisionless) form at lowest collisionality $\nu_{\ast}=10^{-6}$ appear to be split in two independent layers at the intermediate collisionalities $\nu_{\ast}\leq 10^{-4}$ where mismatch of two local maxima located near the middle of field line with the global maximum becomes small compared to the typical boundary layer width (see the discussion of few global maxima case after Eq. (51)). Thus, off-set phenomenon has rather high complexity even in the simple case of a closed field line which is, nevertheless, illustrative for the realistic configurations. In particular, switching off-sets can be observed in Fig. 3 of (Kernbichler et al., 2016) where bootstrap coefficient has been computed for W-7X down to collisionalities $\nu_{\ast}=10^{-9}$.
@@ 3.1 Boundary layer e-folding
As we have seen from asymptotic solutions and numerical examples in both, direct and adjoint problem, distribution function $g_{0}$ which is independent of collisionality in most phase space includes huge contributions localized in boundary layers where they scale as $l_{c}^{1/2}\sim\nu^{-1/2}$ and formally become infinite in the collisionless limit manifested by $\delta$-functions in the asymptotic solutions. To estimate the decay of these localized contributions outside the boundary layer, we consider the homogeneous kinetic equation (3) and use the ansatz of Helander et al. (2011). Namely, we replace in the trapped particle domain the independent variable $\varphi$ with
@@
where the integral $I_{j}=I_{j}(\eta)$ is defined in Eq. (26). Thus, particles with $\lambda>0$ are described by $0<\theta_{H}<\pi$, particles with $\lambda<0$ are described by $-\pi<\theta_{H}<0$, and continuity of the distribution function at both turning points is enabled by periodicity of $g$ with $\theta_{H}$. The homogeneous equation (3) then takes the form
@@
where
@@
Eq. (61) turns into a coupled ODE set for Fourier amplitudes $g_{[m]}(\eta)$. It should be noted that terms with $\partial\theta_{H}/\partial\eta$ which lead to the coupling of Fourier modes are negligible small in most of phase space except for the close vicinity of the boundary (sub-layer (Helander et al., 2011)) where due to the scaling $\partial^{2}\theta_{H}/\partial\eta^{2}\sim\eta^{-1}(\eta-\eta_{b})^{-1}$ their contribution is compatible with the left-hand side which means the sub-layer width $\eta-\eta_{b}\sim\nu_{\ast}\ll\sqrt{\nu_{\ast}}$. Thus, one can ignore the right-hand side terms with derivatives over $\theta_{H}$ outside the sub-layer, and the set decouples into
@@
is actually negligible for the boundary layer contribution because $\delta$-like terms in the direct problem are odd, and in the adjoint problem it describes a jump (55) of the asymptotic solution outside the layer which does not scale with collisionality (this jump is zero in our examples). Ignoring in (64) variation of $D_{H}$ in the main boundary layer, $D_{H}(\eta)\approx D_{H}(\eta_{b})$, we look for solutions in the form $g_{[m]}\propto\exp(\kappa_{m}(\eta-\eta_{b}))$ with $m=1$ harmonic being dominant with increasing $\eta-\eta_{b}$. Thus, boundary layer contribution decays outside the layer as
$g_{0}\approx a\exp\left(\frac{\eta_{b}-\eta}{\delta\eta_{e}}\right)\left(\cos\left(\frac{\eta_{b}-\eta}{\delta\eta_{e}}+\phi\right)\cos\left(|\theta_{H}|\right)-\sigma\sin\left(\frac{\eta_{b}-\eta}{\delta\eta_{e}}+\phi\right)\sin\left(|\theta_{H}|\right)\right),$ (66)
@@
Differences between the numerical and analytical solution of the adjoint problem (boundary layer contribution $\Delta g$) and the asymptotic solution (66) with amplitude $a$ and phase $\phi$ fitted to match $\Delta g$ in the upper left plot are shown in Fig. 9.
@@
Figure 9: Difference $\Delta g=g-g_{0}$ between the result of NEO-2 (the same as in Fig. 6 for $\nu_{\ast}=10^{-4}$) and the analytical solution $g_{0}$ for the adjoint problem (46) (blue), and fit (66) to $\Delta g$ in the vicinity of trapped-passing boundary $\eta_{b}$. Upper left - even and odd parts as functions of $\varphi$ for $\eta-\eta_{b}=0.07$. Upper right - even parts as functions of $\eta$ for $\varphi=19\pi/3$ (7-th local maximum). Lower panel - zooms of the upper right plot over Y-axis.
Oscillations of $\Delta g$ result from interchanging dominance of “particles’ and “anti-particles” which enter the trapping domain and become mirrored there once (twice, etc.). Since the observation point in the last three plots is in the region $\varphi_{b}<\varphi<\varphi_{N}$, “particles” dominate there over “anti-particles” in the first peak (see also Fig. 6) which they produce before the first reflection. Note that similar oscillatory behavior is seen also in the class-transition boundary layers (see Fig. 5) and is reflected then in the small bootstrap off-set oscillation in Fig. 2 at low collisionality where it decreases exponentially.
@@ 3.2 Propagator method, leading order solution
In the following, we use a formal solution of Eq. (69) in terms of its Green’s function, which makes it evident that the last term in Eq. (71) does not contribute to the solution since $\eta$-integral of this term weighted with any function with finite derivative over $\eta$ is zero. The remaining term, as we check later, provides a correction of the order $\delta\eta_{e}$ to the solution driven by $s_{\text{bou}}$, Eq. (70). Thus, we ignore $\Delta s$ fully in the leading order solution. Note that $g_{\text{bou}}$ is an aperiodic function in the passing region due to aperiodicity of $g_{o}^{t}$ in the definition (68). The leading order solution is treated nevertheless as periodic, with the aperiodic term included in the ignored correction driven by $\Delta s$.
We make the following simplifications. As follows from the analysis in Section 3.1, solutions driven by sources located in the passing region or at the trapped-passing boundary tend exponentially to constants of $\eta$ at some sufficiently large distance to the trapped-passing boundary, $\eta>\eta_{m}$, where matching boundary $\eta_{m}$ satisfies $\delta\eta_{e}\ll\eta_{m}-\eta_{b}\ll\eta_{b}$. Generally, these constants can be different in different ripple wells in case they are bounded on both sides by the maxima fulfilling $\eta_{m}B(\varphi_{j})>1$, see Figs. 6 and 7. Thus, we solve Eq. (69) only in the domain $\eta>\eta_{m}$ imposing at the matching boundary zero flux condition, $\partial g_{\text{bou}}(\varphi,\eta_{m},\sigma)/\partial\eta_{m}=0$. The boundary $\eta=\eta_{m}$ serves then as an upper boundary in most trapped particle domain except the interface regions near global maximum and local maxima fulfilling $\eta_{m}B(\varphi_{j})>1$ where $\eta_{m}B(\varphi)>0$, see Fig. 10. Since widths of these regions $\delta\varphi$ are small at low collisionalities, $\delta\varphi\sim B\delta\eta_{e}^{1/2}\left(\partial^{2}B/\partial\varphi^{2}\right)^{-1/2}\propto l_{c}^{-1/4}$, we ignore the collision and source terms there, which leads in such a leading order solution to an error of the order $|\lambda|\delta\varphi\sim(\delta\eta_{e}B)^{1/2}\delta\varphi\sim\delta\eta_{e}B^{3/2}\left(\partial^{2}B/\partial\varphi^{2}\right)^{-1/2}\propto l_{c}^{-1/2}$, being of the same order with an error introduced by ignoring the sub-layer in Section 3.1.
@@
Solution is trivial in the interface regions, and we can express it via mapping between the boundaries $\varphi_{j}\pm\delta\varphi$ where $\varphi_{j}$ is maximum point,
@@
Similar to Section 3.1, we ignore the dependence of diffusion coefficient on $\eta$ setting $D_{\eta}(\varphi,\eta)\rightarrow D_{\eta}(\varphi,\eta_{b})$ and, finally, ignore the extent of interface regions, $\delta\varphi\rightarrow 0$, with each of these steps introducing an error of the order $B\delta\eta_{c}\sim l_{c}^{-1/2}$, as before. Thus, we have transformed our computation domain to a set of rectangular domains $0<\eta<\eta_{m}$, $\varphi_{j}<\varphi<\varphi_{j+1}$, limited by relevant field maxima, $\eta_{m}B(\varphi_{j})>1$ (see Fig. 10). This transformation leads to similar simplifications of the kinetic equation as the transformation of Helander et al. (2011) used in Section 3.1 for the main trapping region (or for the whole field line in case of vanishing boundary layer width) and extended here to multiple local trapping regions.
@@
and initial condition $G_{\sigma}(\varphi^{\prime},\eta,\varphi^{\prime},\eta^{\prime})=\delta(\eta-\eta^{\prime})$, we can formally express the leading order solution to Eq. (69) within a single trapping domain, $\varphi_{j}<\varphi,\varphi^{\prime}<\varphi_{j+1}$, as
@@
and $I_{j}(\eta)$ is given by (26) so that $\delta\eta_{j}=\pi^{1/2}\delta\eta_{e}$, see (67). Conditions (3.2) with $\delta\eta\rightarrow 0$ link the incoming distributions $g^{\pm}_{j(in)}$ to the outgoing $g^{\pm}_{j(out)}$ as follows

### v3

@@ 2.3 Finite collisionality, numerical examples
An even part of the distribution function perturbation $g^{\text{even}}$ responsible for particle flux (Ware pinch) in the adjoint problem is shown in Fig. 6 for the same cases as in Fig. 3. As expected, boundary layer appears at low collisionality only at the trapped-passing boundary. The structure of this layer well agrees with the complete analytical solution (51), i.e. $g^{\text{even}}$ is an anti-symmetric function of $\varphi$ with respect to the middle stellarator symmetry point $\varphi_{b}$.
@@
Figure 7: Even part of the distribution function $g^{\text{even}}$ as function of $\eta$ in the middle of the left off-set well, $\varphi=4\pi/3$ for various collisionalities (left) and $\iota=1/4$. The trapped-passing boundary is shown by a vertical dotted line. The value of $g^{\text{even}}$ at the off-set well bottom, $g_{\text{bot}}$, is shown as a function of $\nu_{\ast}$ in the right plot.
Similar to the direct problem, if the trapped-passing boundary layer is well separated from the nearest class transition boundary (lowest collisionality case, $\nu_{\ast}=10^{-6}$) the Ware pinch coefficient agrees with the asymptotic limit. At higher collisionalities where this boundary layer enters a class-transition boundary, one can observe increased $g^{\text{even}}$ in the local ripple wells whose lowest maximum determines this transition boundary. This accumulation of the particles in local ripples drives the off-set of Ware pinch since $g^{\text{even}}$ there is anti-symmetric with respect to mid-point $\varphi_{b}=5\pi$, what is true also for geodesic curvature and, respectively, for bounce-averaged radial drift velocity in these local ripples. Therefore, particles in both ripples drive the flux of the same sign. In the following, we will call such ripples “off-set wells”. As one can see from Fig. 7, $g^{\text{even}}$ is constant in most of the off-set well, and changes only in the vicinity of class transition boundary. Therefore, we can characterize $g^{\text{even}}$ in the whole off-set well by its value at the off-set well bottom, which will be denoted as $g_{\text{off}}$. It can be seen that $g_{\text{off}}$ has similar feature to $\lambda_{bB}$, i.e. it changes sign with reducing collisionality before going to zero.
@@
Properties of this distribution function off-set, which was also observed in GSRAKE modelling (Beidler, 2020), will be studied in more detail in the next section. Here, we only show an extra example in Fig. 8 where more than one type of off-set wells is present simultaneously with different types of off-sets dominating at different collisionalities. In this example, both, off-set wells containing only the lowest class of trapped particles who traverse a single minimum during their bounce period (as in Fig. 6) and an off-set wells containing many classes are shown. Note that trapped-passing boundary layer which has a standard (collisionless) form at lowest collisionality $\nu_{\ast}=10^{-6}$ appears to be split in two independent layers at the intermediate collisionalities $\nu_{\ast}\leq 10^{-4}$ where mismatch of two local maxima located near the middle of field line with the global maximum becomes small compared to the typical boundary layer width (see the discussion of few global maxima case after Eq. (51)). Thus, off-set phenomenon has rather high complexity even in the simple case of a closed field line which is, nevertheless, illustrative for the realistic configurations. In particular, switching off-sets can be observed in Fig. 3 of (Kernbichler et al., 2016) where bootstrap coefficient has been computed for W-7X down to collisionalities $\nu_{\ast}=10^{-9}$.
@@ 3.1 Boundary layer e-folding
As we have seen from asymptotic solutions and numerical examples in both, direct and adjoint problem, distribution function $g_{0}$ which is independent of collisionality in most phase space includes huge contributions localized in boundary layers where they scale as $l_{c}^{1/2}\sim\nu^{-1/2}$ and formally become infinite in the collisionless limit manifested by $\delta$-functions in the asymptotic solutions. To estimate the decay of these localized contributions outside the boundary layer, we consider the homogeneous kinetic equation (3) and use the ansatz which is only slightly different from the ansatz of Helander et al. (2011) and leads to the same approximation at the end. Namely, we replace in the trapped particle domain the independent variable $\varphi$ with
@@
where the integral $I_{j}=I_{j}(\eta)$ is defined in Eq. (26). Thus, particles with $\lambda>0$ are described by $0<\theta_{H}<\pi$, particles with $\lambda<0$ are described by $-\pi<\theta_{H}<0$, and continuity of the distribution function at both turning points is enabled by periodicity of $g$ with $\theta_{H}$. Formally presenting the homogeneous equation (3) in tensor form
$\frac{\partial}{\partial z^{i}}{\cal J}\left({\cal V}^{i}g-{\cal D}^{ij}\frac{\partial g}{\partial z^{j}}\right)=0,$
where $z^{i}=(\varphi,\eta)$, and transforming the contra-variant components of the phase space velocity ${\cal V}^{i}$ and diffusion tensor ${\cal D}^{ij}$ and the Jacobian $\cal J$ to the new phase space coordinates $(\theta_{H},\eta)$ using tensor algebra rules, this equation takes the form
@@
where differential operators $\partial/\partial\eta$ and $\partial/\partial\theta_{H}$ act on everything to the right, and
@@
Eq. (57) turns into a coupled ODE set for Fourier amplitudes $g_{[m]}(\eta)$. It should be noted that terms with $\partial\theta_{H}/\partial\eta$ which lead to the coupling of Fourier modes are negligible small in most of phase space except for the close vicinity of the boundary (sub-layer (Helander et al., 2011)) where due to the scaling $\partial^{2}\theta_{H}/\partial\eta^{2}\sim\eta^{-1}(\eta-\eta_{b})^{-1}$ their contribution is comparable with the left-hand side which means the sub-layer width $\eta-\eta_{b}\sim\nu_{\ast}\ll\sqrt{\nu_{\ast}}$.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 9: Difference $\Delta g$ and the fit $\Delta g_{\text{fit}}$, Eq. (62) for the case $\nu_{\ast}=10^{-4}$ in Fig. 6. Upper left - even and odd parts as functions of $\varphi$ for $\eta-\eta_{b}=0.07$. Upper right - even parts as functions of $\eta$ for $\varphi=19\pi/3$ (7-th local maximum). Lower panel - zooms of the upper right plot over Y-axis. Dashed line in the last zoom corresponds to NEO-2 result for $g^{\text{even}}-g_{\lambda=0}$ where $g_{\lambda=0}$ is $g^{\text{even}}$ for standing particles with $\eta=1/B$. Function $g_{\lambda=0}$ differs here from $g_{0}=g_{0}^{t}$ by an exponentially small off-set.
Thus, one can ignore the right-hand side terms with derivatives over $\theta_{H}$ outside the sub-layer, and the set decouples into
@@
is actually zero in our case because $\delta$-like terms in the direct problem are odd, and the solution of the adjoint problem is stellarator-antisymmetric. Ignoring in (60) variation of $D_{H}$ in the main boundary layer, $D_{H}(\eta)\approx D_{H}(\eta_{b})$, which brings this equation in agreement with the result of the “rectangular well” ansatz of Helander et al. (2011), we look for solutions in the form $g_{[m]}\propto\exp(\kappa_{m}(\eta-\eta_{b}))$ with $m=1$ harmonic being dominant with increasing $\eta-\eta_{b}$. Thus, boundary layer contribution decays outside the layer as
$\Delta g_{\text{fit}}\approx a\exp\left(\frac{\eta_{b}-\eta}{\delta\eta_{e}}\right)\left(\cos\left(\frac{\eta_{b}-\eta}{\delta\eta_{e}}+\phi\right)\cos\left(|\theta_{H}|\right)-\sigma\sin\left(\frac{\eta_{b}-\eta}{\delta\eta_{e}}+\phi\right)\sin\left(|\theta_{H}|\right)\right),$ (62)
@@
Figure 10: Difference $\Delta g$ (left) and its fit $\Delta g_{\text{fit}}$ (right) as functions of $(\varphi,\eta)$ and $\sigma=1$ (upper panel) and $(\varphi,\lambda)$ (lower panel) for the same case as in Fig. 9. Saturation of color scale at $|\Delta g|\geq 5000$ occurs in the whole passing particle region. Fit $\Delta g_{\text{fit}}$ is plotted excluding passing particle domains and local trapping domains where $\Delta g_{\text{fit}}$ is not defined. Red dotted line in $\Delta g(\varphi,\lambda)$ plot shows the trapped-passing boundary where transient particles move clockwise.
In Figs. 9 and 10, boundary layer contribution represented by the difference $\Delta g=g-g_{0}$ between the numerical (NEO-2) and analytical (Eq. (46)) solution of the adjoint problem is compared to the asymptotic boundary layer solution $\Delta g_{\text{fit}}$ given by Eq. (62) with amplitude $a$ and phase $\phi$ fitted to match $\Delta g$ in the upper left plot of Fig. 9. Oscillations of $\Delta g$ with $\eta$ in Fig. 9 result from interchanging dominance of “particles’ and “anti-particles” which enter the trapping domain and become mirrored there once (twice, etc.). Since the observation point in the last three plots is in the region $\varphi_{b}<\varphi<\varphi_{N}$, “particles” dominate there over “anti-particles” in the first peak (see also Fig. 6) which they produce before the first reflection. The above mechanism results in helical structures in $g(\varphi,\lambda)$ distribution formed in the trapped particle domain (see Fig. 10). They are produced by the combination of the collisionless phase space flow (which is clockwise in this domain) and the flow driven by the collisional diffusion (directed over $\lambda$ to the trapped particle domain from the co- and counter-passing particle domains, where sources of “particles” and “anti-particles” are located, respectively).
Note that oscillatory behavior similar to that in Fig. 9 is seen also in the class-transition boundary layers in Fig. 5 and is reflected then in the small bootstrap off-set oscillation in Fig. 2 at low collisionality where it decreases exponentially.
@@ 3.2 Propagator method, leading order solution
In the following, we use a formal solution of Eq. (65) in terms of its Green’s function, which makes it evident that the last term in Eq. (67) does not contribute to the solution since $\eta$-integral of this term weighted with any function with finite derivative over $\eta$ is zero. The remaining term, as we check later, provides a correction of the order $\delta\eta_{e}$ to the solution driven by $s_{\text{bou}}$, Eq. (66). Thus, we ignore $\Delta s$ fully in the leading order solution. Note that $g_{\text{bou}}$ is an aperiodic function in the passing region due to aperiodicity of $g_{o}^{t}$ in the definition (64). The leading order solution is treated nevertheless as periodic, with the aperiodic term included in the ignored correction driven by $\Delta s$. It should be noted that correction terms ignored here may become important in a special case where the leading order solution tends to vanish. These cases are discussed below in Sections 3.3 and 4.3.
We make the following simplifications. As follows from the analysis in Section 3.1, solutions driven by sources located in the passing region or at the trapped-passing boundary tend exponentially to constants of $\eta$ at some sufficiently large distance to the trapped-passing boundary, $\eta>\eta_{m}$, where matching boundary $\eta_{m}$ satisfies $\delta\eta_{e}\ll\eta_{m}-\eta_{b}\ll\eta_{b}$. Generally, these constants can be different in different ripple wells in case they are bounded on both sides by the maxima fulfilling $\eta_{m}B(\varphi_{j})>1$, see Figs. 6 and 7. Thus, we solve Eq. (65) only in the domain $\eta<\eta_{m}$ imposing at the matching boundary zero flux condition, $\partial g_{\text{bou}}(\varphi,\eta_{m},\sigma)/\partial\eta_{m}=0$. The boundary $\eta=\eta_{m}$ serves then as an upper boundary in most trapped particle domain except the interface regions near global maximum and local maxima fulfilling $\eta_{m}B(\varphi_{j})>1$ where $\eta_{m}B(\varphi)>1$, see Fig. 11. Since widths of these regions $\delta\varphi$ are small at low collisionalities, $\delta\varphi\sim B\delta\eta_{e}^{1/2}\left(\partial^{2}B/\partial\varphi^{2}\right)^{-1/2}\propto l_{c}^{-1/4}$, we ignore the collision and source terms there, which leads in such a leading order solution to an error of the order $|\lambda|\delta\varphi\sim(\delta\eta_{e}B)^{1/2}\delta\varphi\sim\delta\eta_{e}B^{3/2}\left(\partial^{2}B/\partial\varphi^{2}\right)^{-1/2}\propto l_{c}^{-1/2}$, being of the same order with an error introduced by ignoring the sub-layer in Section 3.1. (For this error estimate in the main boundary layer, $|\eta-\eta_{b}|\sim\delta\eta_{e}$, we compared the variance $2\int{\rm d}\varphi D_{\eta}$ in particle $\eta$ introduced by the collisions when traversing the interface region with such variance for the whole trapping region assuming in the latter for simplicity that both, $\lambda$ and region extent in $\varphi$ are of the order one. Error of the above model is of the order one within the sub-layers where effect of collisions on $\lambda$ becomes comparable with the effect of mirroring force, but the overall contribution of these narrow sub-layers is similarly small.)
@@
The solution is trivial in the interface regions, and we can express it via mapping between the boundaries $\varphi_{j}\pm\delta\varphi$ where $\varphi_{j}$ is maximum point,
@@
Similar to Section 3.1, we ignore the dependence of diffusion coefficient on $\eta$ setting $D_{\eta}(\varphi,\eta)\rightarrow D_{\eta}(\varphi,\eta_{b})$ and, finally, ignore the extent of interface regions, $\delta\varphi\rightarrow 0$, with each of these steps introducing an error of the order $B\delta\eta_{c}\sim l_{c}^{-1/2}$, as before. Thus, we have transformed our computation domain to a set of rectangular domains $0<\eta<\eta_{m}$, $\varphi_{j}<\varphi<\varphi_{j+1}$, limited by relevant field maxima, $\eta_{m}B(\varphi_{j})>1$ (see Fig. 11). This transformation leads to similar simplifications of the kinetic equation as the transformation used in Section 3.1 for the main trapping region (or for the whole field line in case of vanishing boundary layer width) and is the same as the transformation of Helander et al. (2011) extended here to multiple local trapping regions.
@@
and initial condition $G_{\sigma}(\varphi^{\prime},\eta,\varphi^{\prime},\eta^{\prime})=\delta(\eta-\eta^{\prime})$ at the starting point $\varphi^{\prime}$, we can formally express the leading order solution to Eq. (65) within a single trapping domain, $\varphi_{j}<\varphi,\varphi^{\prime}<\varphi_{j+1}$, as
@@
and $I_{j}(\eta)$ is given by (26) so that $\delta\eta_{j}=\pi^{1/2}\delta\eta_{e}$, see (63). Conditions (3.2) with $\delta\varphi\rightarrow 0$ link the incoming distributions $g^{\pm}_{j(in)}$ to the outgoing $g^{\pm}_{j(out)}$ as follows